OpenStudy (anonymous):

1. remainder when 2^91 is divided by 7 ? 2. remainder when 25! is divided by 10^7 ? please explain the approach :) thank you !

OpenStudy (anonymous):

\(2^2=2,2^2=4,2^3\) patter will repeat remainder when divided by 7 of each of these is \(2,4,1\)

OpenStudy (anonymous):

*pattern

OpenStudy (anonymous):

ok.. understood so far ! :)

OpenStudy (anonymous):

we can check the next ones \(2^4=16\) \(2^5=32, 2^6=64\) remainders are again \(2,4,1\)

OpenStudy (anonymous):

so repeats every 3 now \(2^{91}\) should be ok right?

OpenStudy (anonymous):

answer is 2 ?

OpenStudy (anonymous):

i think so yes

OpenStudy (anonymous):

thats brilliant :D thank you!

OpenStudy (anonymous):

since 9 goes in to 90 evenly, the remainder for \(2^{90}=1\) and so remainder of \(2^{91}\) would be 2

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

as for the next one i am going to guess at 0 but we need to know how many zeros are in \(25!\)

OpenStudy (anonymous):

no that is wrong, damn

OpenStudy (anonymous):

because there are only 6 zeros in \(25!\) so you need the last non zero digit i am not sure how to do this without a direct computation though

OpenStudy (anonymous):

the answer is 4 but that is by computing. i can't think of a snap way to do it let me know if you come up with one

OpenStudy (anonymous):

sure :)